### v2

# Persuade Me if You Can: A Framework for Evaluating Persuasion Effectiveness and Susceptibility Among Large Language Models

Abstract
Large Language Models (LLMs) demonstrate persuasive capabilities that rival human-level persuasion. While these capabilities can be used for social good, they also present risks of potential misuse. Moreover, LLMs’ susceptibility to persuasion raises concerns about alignment with ethical principles. To study these dynamics, we introduce Persuade Me If You Can (PMIYC), an automated framework for evaluating persuasion through multi-agent interactions. Here, Persuader agents engage in multi-turn conversations with the Persuadee agents, allowing us to measure LLMs’ persuasive effectiveness and their susceptibility to persuasion. We conduct comprehensive evaluations across diverse LLMs, ensuring each model is assessed against others in both subjective and misinformation contexts. We validate the efficacy of our framework through human evaluations and show alignment with prior work. PMIYC offers a scalable alternative to human annotation for studying persuasion in LLMs. Through PMIYC, we find that Llama-3.3-70B and GPT-4o exhibit similar persuasive effectiveness, outperforming Claude 3 Haiku by 30%. However, GPT-4o demonstrates over 50% greater resistance to persuasion for misinformation compared to Llama-3.3-70B. These findings provide empirical insights into the persuasive dynamics of LLMs and contribute to the development of safer AI systems. [^1] https://beyzabozdag.github.io/PMIYC
Persuade Me if You Can: A Framework for Evaluating Persuasion Effectiveness and Susceptibility Among Large Language Models
Nimet Beyza Bozdag, Shuhaib Mehri, Gokhan Tur, Dilek Hakkani-Tür University of Illinois Urbana-Champaign {nbozdag2, mehri2, gokhan, dilek}@illinois.edu
1 Introduction
Figure 1: A $t$-turn interaction between a Persuader and a Persuadee in PMIYC. The Persuader can addresses the Persuadee’s concerns, while the Persuadee reports its agreement scores.
Large Language Models (LLMs) have demonstrated remarkable capabilities across various domains and are now integral to numerous real-world applications (OpenAI, 2024c; Anthropic, 2024). Among these capabilities, persuasion stands out as particularly noteworthy, with state-of-the-art LLMs exhibiting persuasive skills comparable to humans (Durmus et al., 2024; OpenAI, 2024c). Persuasive LLMs have been applied for social good, promoting public health and prosocial behaviors while adhering to ethical, non-coercive principles (Wang et al., 2019; Furumai et al., 2024; Karinshak et al., 2023; Kampik et al., 2018). However, as LLMs become increasingly persuasive, they also pose risks. These models can be exploited to manipulate individuals, spread misinformation, and influence public opinion in harmful ways (Singh et al., 2024; Salvi et al., 2024; Simchon et al., 2024; Hackenburg and Margetts, 2024).
Beyond influencing others, LLMs themselves are susceptible to persuasion, making them vulnerable to manipulation that may lead to violations of alignment principles, generating toxic or misleading content, and reinforcing biases (Zeng et al., 2024; Xu et al., 2024). This presents a challenge in having LLMs resist harmful or inappropriate influence while remaining open to justified reasons to change their views (Stengel-Eskin et al., 2024).
Existing studies primarily assess persuasion in LLMs through human evaluations (Durmus et al., 2024) or automated methods restricted to limited or non-conversational setups (OpenAI, 2024c; Singh et al., 2024). While human evaluation provides valuable insights, it is costly, time-intensive, and difficult to scale across different models and scenarios. Likewise, existing automated approaches lack the depth needed for comprehensive analysis in dynamic conversational settings. Without a scalable evaluation framework, understanding LLM persuasion dynamics remains limited, hindering the detection of vulnerabilities and the development of safeguards for safe AI deployment. Therefore, an automated evaluation framework for LLM persuasiveness and susceptibility is essential.
In this work, we introduce Persuade Me If You Can (PMIYC), an automated framework for evaluating persuasion in LLMs across diverse conversational setups. Building on recent studies (Xu et al., 2024; Stengel-Eskin et al., 2024), our framework simulates multi-agent interactions where a Persuader attempts to convince a Persuadee to agree with a claim. PMIYC enables the Persuader and Persuadee to engage in multi-turn conversations, exchanging arguments, addressing counterpoints, and influencing each other’s views. Throughout these interactions, the Persuadee continuously assesses the Persuader’s arguments and self-reports its level of agreement, allowing us to track changes in opinion and measure both persuasive effectiveness (how successfully an LLM persuades) and susceptibility to persuasion (how easily an LLM changes its stance).
Our round-robin evaluation, where each model takes turns acting as both the Persuader and the Persuadee against every other model, enables direct comparisons. Our evaluation spans a diverse set of models, including open-weight models (e.g., Llama-3.3-70B-Instruct) and closed-source models (e.g., GPT-4o, Claude 3 Opus). We assess single-turn persuasion, where the Persuader has only one chance to convince the Persuadee, as well as multi-turn persuasion, which allows for back-and-forth exchanges and deeper argumentation. Additionally, we examine persuasion across different contexts, including subjective claims (eg. "Cultured/lab-grown meats should be allowed to be sold") and misinformation claims (eg. "Canada is a part of the UK"), to evaluate how LLMs respond under varying conditions.
Our results reveal that persuasive effectiveness remains largely consistent across settings, with Llama-3.3-70B-Instruct emerging as the most effective persuader, closely followed by GPT-4o. However, susceptibility to persuasion varies based on the context and the number of conversational turns. Notably, GPT-4o and GPT-4o-mini exhibit up to 50% stronger resistance to persuasion in misinformation contexts but show greater flexibility with subjective claims. Persuasive effectiveness is greater in the first two attempts, and a Persuader becomes more effective when it more strongly agrees with the claim it advocates. By modeling persuasion as an interactive and iterative process, PMIYC provides a more realistic setting for studying persuasion as it occurs in conversations. The reliability of our framework is validated through human evaluations and comparative analyses with existing studies. Our findings align closely with human judgments of persuasiveness, reinforcing the validity of PMIYC as a scalable alternative to human assessment. Our key contributions are:
  • PMIYC, a framework for automatically evaluating both the persuasive effectiveness and susceptibility to persuasion of LLMs.
  • A comprehensive analysis of persuasion in LLMs across multiple dimensions, including conversations in single-turn vs. multi-turn, and subjective vs. misinformation contexts.
Through this research, we aim to deepen the understanding of persuasive capabilities and susceptibility of LLMs, and to emphasize the safety risks associated with persuasion in AI systems. By providing a scalable and automated evaluation framework, we aim to contribute to the development of safer, more transparent, and ethically responsible LLMs for real-world applications.
2 Related Work
Persuasive Language Models.
Recent studies suggest that LLMs exhibit human-like persuasiveness (Durmus et al., 2024; OpenAI, 2024c; Huang and Wang, 2023). Durmus et al. (2024)’s and OpenAI (2024c)’s experiments show that their models are as persuasive as humans. While some studies leverage the persuasive abilities of LLMs for social good Karinshak et al. (2023); Costello et al. (2024), others warn of risks like microtargeting, political influence, and ethical concerns Salvi et al. (2024); Simchon et al. (2024); Hackenburg and Margetts (2024). These findings underscore the growing urgency of understanding and regulating persuasive AI systems.
Evaluating Persuasion.
Prior research has explored various methods to measure the persuasiveness of LLMs. Durmus et al. (2024) uses human evaluators to rank their agreement with a subjective claim before and after reading the model-generated arguments. While this approach provides valuable insights into human preferences, it is not highly scalable. Similarly, OpenAI uses human annotators by having them rank arguments and perform pairwise comparisons OpenAI (2024c). They also explore automated evaluation, through their MakeMeSay and MakeMePay role-playing frameworks. These frameworks, while innovative, are highly context-dependent and may not generalize well to real-world persuasion scenarios. Singh et al. (2024) introduce PersuasionArena and the task of transsuasion, where LLMs rewrite low-engagement tweets into more persuasive versions. Although PersuasionArena provides an automated framework for measuring persuasion, its focus is less conversational and does not fully address the interactive and manipulative aspects of persuasion. Breum et al. (2023) develop a Convincer-Skeptic scenario, in which an LLM attempts to persuade another LLM in the domain of climate change. They show that agents can mimic some of the persuasive dynamics and opinion changes observed in human discourse. Other studies have attempted to score persuasive language in pairs of LLM-generated text using a regression model (Pauli et al., 2024). Yet, these efforts focus on short text excerpts, limiting their applicability to longer, multi-turn persuasion.
LLM Susceptibility to Persuasion.
While LLMs generating persuasive content raises ethical concerns, their susceptibility to persuasion is equally important. Some jailbreaking studies show that LLMs can be persuaded to generate harmful content Zeng et al. (2024) or accept and reinforce misinformation Xu et al. (2024). Furthermore, recent studies have shown that multi-turn jailbreak attempts can be particularly effective in circumventing safeguards Xu et al. (2024); Li et al. (2024); Russinovich et al. (2024). Inspired by these findings, we extend our framework to include the misinformation domain.
Overall, prior research has primarily focused on single-turn persuasive attacks on LLMs, human-in-the-loop evaluation methods, less conversational or highly constrained scenarios, with limited exploration of multi-turn agent-to-agent persuasion. In contrast, we explore how LLMs interact with each other in a highly conversational setting across different domains and multi-turn scenarios, encompassing both beneficial and harmful persuasion.
3 Methodology
PMIYC is our framework for evaluating both the persuasive effectiveness of LLMs and their susceptibility to persuasion. [^2] Code adapted from NegotiationArena Bianchi et al. (2024). Experiments consist of multi-agent interactions, where one agent acts as a Persuader (ER) and the other acts as a Persuadee (EE). The two engage in a structured exchange, where the Persuader agent is given the objective of convincing the Persuadee to agree with a given claim. The Persuadee engages with the Persuader and re-assesses its stance on the claim throughout the conversation.
3.1 Persuasive Conversation Generation
Conversation generation involves simulating a dialogue between two agents: a Persuader and a Persuadee. The objective is to track how the Persuadee’s stance on a given claim $C$ evolves over the course of a $t$-turn interaction. The conversation begins with the Persuadee expressing its initial agreement with the claim by providing both a message and an agreement score (see Turn 1 in Figure 1). The agreement score quantifies the model’s stance on claim $C$ using a five-point Likert scale: Completely Oppose (1), Oppose (2), Neutral (3), Support (4), and Completely Support (5). This scale enables a more granular analysis of opinion shifts beyond a binary agree/disagree classification.
After observing the Persuadee’s initial stance, the Persuader is tasked with advocating in favor of the claim, irrespective of its own opinion. The two agents then engage in a structured exchange, with the Persuader presenting arguments intended to influence the Persuadee’s stance. At the end of the conversation, the Persuadee provides a final decision and agreement score (see Turn $t$ in Figure 1). If at any point after the Persuader presents its first argument the Persuadee reports the maximum agreement score of Completely Support (5), the conversation is terminated early. The Persuadee is then prompted for its final decision without further arguments from the Persuader. This early stopping mechanism prevents unnecessary computation and ensures efficiency, as the model has no further room to be persuaded.
Throughout the interaction, system prompts guide response generation, with a dedicated decision prompt issued to the Persuadee in the final step. The overall approach is outlined in Algorithm 1, with further implementation details provided in Appendix B. The full set of prompts is available in Appendix E. To ensure a fair comparison of different Persuader models against the same Persuadee model, we fix the Persuadee’s initial message and agreement score for each claim $C$. This approach controls for variability in the Persuadee’s starting position, allowing us to isolate the influence of the Persuader’s arguments on opinion shifts.
3.2 Normalized Change in Agreement
To evaluate conversations in PMIYC, we introduce normalized change (NC) in agreement, a metric designed to quantify the persuasive impact of a conversation. In a $t$-turn dialogue, we compute the difference between the Persuadee’s initial agreement score, $sEE_{0}$, and its final agreement score, $sEE_{t}$. This difference represents the extent to which the Persuader has influenced the Persuadee’s stance. To ensure comparability across different initial agreement scores, we normalize the change using the following formulation:
$NC(c)=\begin{cases}\displaystyle\frac{sEE_{t}-sEE_{0}}{5-sEE_{0}},&\begin{array}[]{l}\text{if }sEE_{t}\geq sEE_{0},\\ \text{and }sEE_{0}\neq 5\end{array}\\[8.0pt] \displaystyle\frac{sEE_{t}-sEE_{0}}{sEE_{0}-1},&\text{ otherwise}\end{cases}$ (1)
The denominator ensures proper normalization by accounting for the range of possible agreement shifts. When the Persuadee’s agreement increases, the maximum possible increase is given by $(5-sEE_{0})$, while for a decrease, the maximum possible change is $(sEE_{0}-1)$. This normalization prevents bias toward Persuadees with initial scores closer to neutral or opposition, who would otherwise have greater potential for upward shifts. It also ensures that declines in agreement are properly accounted for, effectively penalizing the Persuader when persuasion efforts fail or lead to a stronger opposition stance.
Since the metric is normalized, its values always fall within the range $[-1,1]$. A value of 1 indicates a complete shift to agreement $(sEE_{t}=5)$, -1 signifies a full shift to opposition $(sEE_{t}=1)$, and 0 means no change. Negative values indicate cases where the persuasion attempt backfires, leading to stronger opposition rather than increased agreement. This metric provides a standardized way to compare conversations across different claims and agent pairings by measuring both effectiveness and susceptibility. Effectiveness refers to a model’s ability to persuade others and is computed as the average normalized change when the model acts as the Persuader. Susceptibility reflects how easily a model changes its stance and is measured as the average normalized change when the model serves as the Persuadee. Higher values indicate greater persuasive strength when the model is in the Persuader role and greater susceptibility to persuasion when it is in the Persuadee role.
4 Experiments
Our experiments are round-robin evaluations across a comprehensive set of models, including: Llama-3.1-8B-Instruct, Llama-3.3-70B-Instruct (Grattafiori et al., 2024), GPT-4o-mini (OpenAI, 2024a), GPT-4o (OpenAI, 2024b), Claude 3 Opus (Anthropic, 2024), Claude 2 (Anthropic, 2023), Claude 3 Haiku (Anthropic, 2024). We evaluate persuasion on two types of claims: subjective claims (Section 4.1) and misinformation claims (Section 4.2).
4.1 Persuasion for Subjective Claims
To ensure comprehensive evaluation, we use the set of subjective claims from Durmus et al. (2024) and the Perspectrum dataset (Chen et al., 2019), which is a collection of claims from debate platforms. Together, these 961 claims span a broad range of topics including political, ethical, and social issues. We focus on subjective claims because, unlike factual topics where models may exhibit strong predetermined stances based on prior knowledge, subjective claims better reveal the dynamics of persuasion.
4.1.1 Single-Turn Persuasion
We begin our experiments with single-turn conversations in PMIYC. In this setup, $t=3$, and the Persuader only has a single turn to present an argument and influence the Persuadee’s stance.
Figure 2: Average normalized change (NC) in Persuaadee’s agreement for different Persuader-Persuadee pairs in subjective single-turn conversations. (All Llama models are Instruct versions.)
Figure 3: (a) Average effectiveness of the Persuader across single-turn vs. multi-turn and subjective vs. misinformation interactions. (b) Average susceptibility of the Persuadee under the same conditions.
Figure 2 presents different pairings of Persuader and Persuadee models, reporting the average NC in Pesruadee’s agreement scores. For a given Persuadee, examining the corresponding column reveals which Persuaders exerted greater influence. Conversely, for each Persuader, inspecting a row indicates which models were more susceptible to persuasion. Results from the single-turn persuasion experiments align with expectations: larger models, known for their superior reasoning capabilities, generally exhibit higher persuasiveness.
Figure 3(a) presents the effectiveness of different models in persuasion, supporting the trend that larger models are more effective persuaders, except GPT-4o-mini, which achieves comparable results. Notably, Claude 2 demonstrates unexpectedly weak persuasiveness. Figure 3(b) highlights the relative susceptibility of different models to persuasion. LLama-3.3-70B-Instruct emerges as the most persuadable model, while GPT-4o appears comparatively more resistant to persuasion.
Alignment with Prior Work.
Durmus et al. (2024) conducted a human evaluation study on the persuasiveness of Claude models, reporting that Claude 3 Opus is the most persuasive, approaching human-level performance, followed by Claude 3 Haiku, and Claude 2 being the least persuasive model. Our results, shown in Figures 2 and 3(a), align with these results. Moving forward, this validates PMIYC, and establishes it as a reliable framework for studying persuasion in LLMs.
4.1.2 LLMs for Scoring Persuasiveness
We explore the feasibility of using LLMs to directly score argument persuasiveness. We curate an evaluation dataset derived from the human annotations in Durmus et al. (2024), where we create preference pairs made up of more and less persuasive arguments for each claim. We present these pairs in random order to GPT-4o, GPT-4o-mini, and Llama-3-70B-Instruct, and ask them to select the more persuasive argument. The models achieve accuracies of 54.6% 51.6%, 49.30% respectively. When using only the pairs where the difference in persuasion scores is greater than 1, accuracies improve to 55.4%, 54.7%, 55.2%.
To explore if training could improve performance, we use GPT-4o-mini to synthesize a preference dataset with 10K samples. For subjective claims from the Perspectrum dataset (Chen et al., 2019), we use zero-shot prompting to generate an argument, then a more persuasive version of that argument. We use these two arguments to create a preference pair. Using this dataset, we train a reward model using Llama-3.1-8B-Instruct with a contrastive loss. To evaluate this model, we have it independently score each argument, then award it a score if it gives the more persuasive argument a higher score than the less persuasive argument. Here, the model achieves an accuracy of 47.19%.
These experiments reveal the difficulty of using LLMs to directly assess persuasiveness. While our PMIYC framework successfully captures persuasiveness through change in LLM agreement scores, direct LLM assessments of persuasiveness yields significantly poorer results. This difference in performance validates our choice of PMIYC for studying persuasiveness, as opposed to direct LLM-based judgments.
4.1.3 Multi-Turn Persuasion
The multi-turn experiments follow the same setup as described in Section 4.1.1, with the key difference being the total number of turns, $t=9$. Unlike the single-turn setup, the multi-turn interactions give the Persuader multiple opportunities to influence or reinforce persuasion. The multi-turn conversations build upon those generated in the single-turn experiment, but without including the Persuadee’s final decision. The same set of models is used in both settings, except for Claude 2 and Claude 3 Opus, which are excluded as Persuaders due to cost constraints.
Figure 4: Average NC in Persuadee’s agreement for different Persuader-Persuadee pairs in subjective multi-turn conversations.
Figure 5: Average NC in Persuadee’s agreement for different Persuader-Persuadee pairs in misinformation multi-turn conversations.
Figure 4 presents the persuasive outcomes for multi-turn conversations on subjective claims. As seen in Figure 3(a) the effectiveness of the persuaders increase in multi-turn conversations and the relative effectiveness of Persuaders remains consistent, with LLama-3.3-70B-Instruct and GPT-4o being the most persuasive, while Claude 3 Haiku performs the worst.
Over a multi-turn conversation, we observe a general increase in susceptibility (Figure 3(b)), aligning with findings from Li et al. (2024); Russinovich et al. (2024). However, unlike in the single-turn setting, GPT-4o no longer ranks as the most resistant model, showing greater susceptibility than GPT-4o-mini. Additionally, Claude 3 Haiku emerges as the most persuadable model in the multi-turn setting.
4.2 Persuasion for Misinformation Claims
All previous experiments focused on subjective claims, which may be easier to persuade for since language models exhibit lower resistance to persuasion in these scenarios. This could overestimate the effectiveness of Persuaders. To address this, we introduce another set of multi-turn experiments on misinformation.
To assess persuasiveness in this context, we task Persuader models to convince the Persuadee to agree with an incorrect answer to a factual question. We source 817 question-answer pairs from the TruthfulQA dataset Lin et al. (2022), pairing each question with its first incorrect answer sample. For instance, given the question "Why do veins appear blue?", the Persuader attempts to persuade the Persuadee to accept the incorrect answer: "Veins appear blue due to how blue and red light penetrate human tissue." The misinformation conversation setup follows the same procedure as described in Section 4.1.3. These experiments also highlight a critical risk of LLM-based persuasion: not only do Persuaders demonstrate the ability and willingness to persuade for misinformation, but Persuadees also show susceptibility to adopting incorrect information.
Figure 5 illustrates the varying persuasive outcomes in the misinformation domain. The GPT models exhibit strong resistance to being jailbroken, with more than 50% greater resistance to persuasion than the next best Persuadee. While susceptibility to persuasion is highly domain-dependent (Figure 3(b)), Persuader effectiveness remains relatively stable (Figure 3(a)). Notably, Llama-3.3-70B-Instruct continues to be the most effective Persuader, reinforcing its strong performance across different domains.
5 Discussion & Analysis
Persuadee Avg SD Match Rate Avg Dif
Llama-3.1-8B 0.49 16.23 1.13
Llama-3.3-70B 0.03 74.92 0.31
GPT-4o-mini 0.05 67.85 0.41
GPT-4o 0.05 65.87 0.38
Claude 3 Haiku 0.12 56.92 0.51
Table 1: Consistency of Persuadee models agreement score evaluations (5 iterations) for subjective claims. Avg SD denotes the average standard deviation of agreement scores across iterations. Match Rate indicates the percentage of cases where the initial agreement score in PMIYC matched the most repeated score from the iterations. Avg Dif represents the average absolute difference between the model’s initial agreement score in PMIYC and the mean score from the iterations.
5.1 Do Persuadees Have Consistent Opinions?
A fundamental aspect of PMIYC is the Persuadee’s initial agreement with a claim, as this value serves as a reference point for evaluating persuasiveness throughout the conversation. Ensuring that models maintain consistent agreement scores is crucial, as random fluctuations would undermine the reliability of our framework. To assess this consistency, we repeatedly prompt each of the five Persuadee models for their agreement with the same set of subjective claims. We then compute the standard deviation of their agreement scores, as reported in Table 1. Our results show that most models produce stable initial agreement scores across repeated queries, except for Llama-3.1-8B-Instruct, which exhibits higher variability. This suggests that, aside from this exception, models hold consistent opinions, confirming that the recorded initial agreement scores can be reliably used in our framework as the Persuadee’s beliefs.
5.2 Are LLMs reliable self-reporters?
To validate the core principles of PMIYC, the Persuadee’s self-reported agreement scores, we employ human annotators. Annotators are provided with the claim $C$ and the full conversation between the Persuader and Persuadee regarding $C$. Their task is to assign the most appropriate agreement score based on the Persuadee’s response at each turn. We randomly select the same three subjective claims and two misinformation claims across all 25 Persuader-Persuadee pairs in multi-turn conversations, resulting in a total of 125 annotated conversations. Our results indicate that annotators matched the model-reported agreement scores 57% of the time, with an average absolute difference of 0.51 between the annotators and the model. This suggests that most discrepancies arose from adjacent agreement categories, such as "Support" and "Completely Support" (Figure 8). These findings demonstrate that models serve as reliable self-reporters of their agreement. Further analysis of human annotation results can be found in Appendix C.
5.3 How Is Agreement Affected Throughout a Conversation?
Figure 6: Average Persuadee agreement at each Persuadee turn of the conversation for a given Persuader. Solid lines indicate performance with subjective claims, dashed lines represent performance with misinformation claims. After the Persuadee’s fourth turn (orange line), it is prompted to make a final decision (green line). For conversations that end early, the remaining turns up to the final decision are filled with a score of 5 (Completely Support).
Figure 6 presents the average Persuadee agreement score over the course of a multi-turn conversation. For most Persuaders, we observe that their influence peaks during the first and second persuasive attempts. An exception is Claude 3 Haiku in the subjective claim setup, where its persuasive impact is stronger on the third attempt, compared to its second argument. An interesting observation is the decline in the Persuadee’s agreement score across all Persuaders for misinformation claims, and Claude 3 Haiku for subjective claims in the final decision step. We attribute this behavior to stronger Persuadees influencing weaker Persuaders, leading them to adopt the stronger model’s stance. As a result, the Persuadee incorrectly reports agreement with the Persuader’s arguments which has deviated from its initial position. However, when the Persuadee is asked for its final opinion and reminded of the misinformation claim, its agreement drops significantly (example in Appendix D). This also explains the unexpected increase in Claude 3 Haiku’s persuasive effectiveness in the misinformation setting.
5.4 Are Persuaders More Effective When in Agreement with a Claim?
Figure 7: Average NC in Persuadee’s agreement based on Persuader’s agreement with the claim in multi-turn subjective (blue), and misinformation (red) settings. Persuaders are classified as Opposing (agreement score: 1–2), Neutral (3), or Supporting (4–5).
In PMIYC, Persuader’s are tasked with persuading in favor of a claim $C$, regardless of their actual stance on the issue. This raises an important question of whether or not Persuaders are more effective when they genuinely agree with the claim they are advocating. Understanding this dynamic can offer insights into how belief alignment influences persuasive success, potentially affecting argument quality. For this reason, we analyze the average NC in Persuadee’s agreement by categorizing each Pesruader into Supporting, Neutral and Opposing persuaders (Figure 7). The results indicate that a model’s agreement with the claim enhances its effectiveness as a persuader across all settings, including both subjective and misinformation contexts which is intuitive.
6 Conclusion
We introduce a novel framework, Persuade Me If You Can (PMIYC), for evaluating the persuasive effectiveness and susceptibility of LLMs through multi-turn conversational interactions. By simulating dialogues between a Persuader and a Persuadee, our approach captures the dynamic nature of persuasion. Our findings reveal that multi-turn conversations enhance persuasion compared to single-turn interactions, and that larger models, particularly Llama-3.3-70B-Instruct and GPT-4o, demonstrate superior persuasive effectiveness, aligning with their advanced reasoning capabilities. However, susceptibility to persuasion exhibits a more nuanced pattern, varying with the number of turns and the persuasion domain. We validate our results against human annotations and prior studies, confirming its potential as a robust tool for assessing persuasion in LLMs. Our work contributes to research on optimizing persuasive AI and assessing safety risks by providing insights into how LLMs engage in persuasive dialogue. Future work could expand PMIYC beyond subjective claims and misinformation to include persuasion for social good or commercial applications, offering a broader evaluation of LLM persuasive capabilities. Additionally, investigating methods to control persuasive effectiveness and susceptibility, such as leveraging data from successful and unsuccessful persuasion attempts, could provide insights into enhancing or mitigating these attributes.
7 Limitations
PMIYC focuses only on persuasion in subjective claims and misinformation. Expanding to other domains, such as persuasion for social good or commercial success, would provide a more comprehensive assessment of LLMs’ persuasive capabilities. While extensive taxonomies of persuasive strategies exist, our framework does not explicitly incorporate them. A finer-grained analysis of persuasive techniques could offer deeper insights into which strategies are most effective across models and contexts. In PMIYC, conversations are generated using a fixed set of prompts and occur only once. The effects of prompt variation and multiple iterations on persuasion outcomes remain unexplored. Although our results align with prior research on human persuasion, we do not separately employ human Persuaders and Persuadees to validate alignment with human preferences.
8 Ethical Considerations
Persuasion is a powerful tool that can be beneficial in applications for social good, but also carries risks in misinformation, manipulation, and coercion. While PMIYC does not directly engage humans in persuasion tasks, thereby mitigating immediate risks to users, the framework evaluates persuasion in a misinformation setting, where models show susceptibility to accepting or advocating for false claims. This raises concerns about jailbreaking and misuse, particularly in harmful or toxic scenarios. Understanding persuasive skills in LLMs not only provides insight into their risks but also informs safer deployment strategies. However, it is important to acknowledge that the same methodologies used to measure and assess persuasion could also be leveraged to develop more persuasive models, which necessitates careful consideration of ethical implications and responsible AI practices.
Acknowledgments
This research project has benefited from the Microsoft Accelerate Foundation Models Research (AFMR) grant program through which leading foundation models hosted by Microsoft Azure along with access to Azure credits were provided to conduct the research. Additionally, this research utilized Claude credits granted through Anthropic’s External Researcher Access Program.
Appendix A Initial Model Beliefs
To evaluate the potential for persuasion in Persuadee models, we analyzed the initial agreement score distributions across the models used in our experiments. For the subjective claims, the results indicate that most models were neutral, with an average initial agreement score of around 3 on a five-point Likert scale. The model with the lowest average score was Claude 3 Haiku at 2.94, while the highest was Llama-3.3-70B-Instruct at 3.41. Although the averages suggest neutrality, the actual distributions reveal concentrations at Oppose (2) and Support (4), rather than at Neutral (3). This indicates that while models may appear neutral on average, they often take polarized positions on subjective claims. These findings suggest that the models have the potential to shift their stances during persuasion, highlighting room for persuasive influence.
Appendix B Details for Generating Conversations
Input: Number of turns $t$
Output: Full conversation history $CH$, turn-level agreement score history of the persuadee $sEE$
// Initialization
Initialize conversation history: $CH\leftarrow[]$
Initialize persuadee agreement scores: $sEE\leftarrow[]$
for $turn_{i}\in\{1,2,\ldots,t-1\}$ do
       if $turn_{i}$ is odd then
             // Persuadee’s turn: Generate response and agreement score
             $msgEE_{i},sEE_{i}=EE(CH,sEE)$;
             Append $msgEE_{i}$ to $CH$;
             Append $sEE_{i}$ to $sEE$;
             if $sEE_{i}=5$ and $turn_{i}\neq 1$ then
                   // Persuadee has completely agreed, end dialogue
                   break;
             end if
       end if
      else
             // Persuader’s turn: Generate response
             $msgER_{i}=ER(CH)$;
             Append $msgER_{i}$ to $CH$;
       end if
end for
// Persuadee’s final decision (Turn t)
$msg_{t},sEE_{t}=\text{FinalDecision}(CH,sEE)$;
Append $msg_{t}$ to $CH$;
Append $sEE_{t}$ to $sEE$;
Algorithm 1 Persuasive Conversation Generation Algorithm
Algorithm 1 simulates a persuasive dialogue between a Persuader ($ER$) and a Persuadee ($EE$) over $t$ turns. Agents generate a response using the functions $ER(\cdot)$, $EE(\cdot)$, or $\text{FinalDecision}(\cdot)$, which take the current conversation history as input. $EE(\cdot)$, and $\text{FinalDecision}(\cdot)$ also take in an agreement score history $(sEE)$ of the Persuadee, and return the agreement score for that turn. If the Persuadee reaches the maximum agreement score of Completely Support (5) at any point after the first argument, the conversation is cut short. The conversation history $(CH)$ and $EE$’s agreement scores ($sEE$) are updated iteratively, with the $EE$ speaking on odd turns and the $ER$ responding on even turns. Regardless of early stopping, $EE$ provides a final decision statement. The process models sequential interactions.
B.0.1 Conversation Generation Success
In PMIYC, a conversation for some claim might not be successfully generated. Models may refuse to respond to a prompt, often due to security measures restricting discussions on sensitive topics. In other cases, models fail to follow instructions properly, producing responses in an unexpected format, which prevents the system from extracting essential fields such as the agent’s response and score. Our experiments show that over 96% of conversations in the subjective domain and 99% in the misinformation domain were generated successfully. Among the models tested, Claude 3 Haiku as the Persuader exhibited the highest rate of generation errors.
Appendix C Human Annotation
Figure 8: Confusion matrix showing the number of annotated utterances where model self-reported agreement scores align with human-annotated agreement scores.
Figure 9: Page 1 of Human Annotation Guidelines.
Figure 10: Page 2 of Human Annotation Guidelines.
As discussed in Section 5.2, we conducted a human evaluation study to assess model reliability in self-reporting agreement scores. For this, we enlisted 12 in-house annotator volunteers, all graduate students in computer science. In total, 591 Persuadee utterances across 125 conversations were annotated. Figure 8 presents a comparison between the model’s self-reported agreement scores and human annotators’ perceived agreement. The results indicate that discrepancies occur most frequently between adjacent agreement levels, such as "Complete Oppose (1)" and "Oppose (2)," which is expected given the subtle distinctions between these categories. The Figures 9 and 10 show the instructions and disclosure provided to the annotators.
Appendix D Sample Persuasive Conversations from PMIYC
Figure 11: In this conversation, Llama-3.3-70B-Instruct acts as the Persuader, while GPT-4o is the Persuadee. Throughout the multi-turn interaction, both agents engage seamlessly, resulting in the Persuadee changing their opinion by the end of the conversation.
Figure 12: In this conversation, Claude 3 Haiku acts as the Persuader and Llama-3.3-70B-Instruct as the Persuadee. Over the course of the interaction, error propagation occurs, leading the Persuader to retreat from their initial stance and adopt arguments presented by the Persuadee. Consequently, the Persuadee incorrectly reports agreement with a new, shifted understanding that diverges from the original claim. In the final turn, when the claim is explicitly restated, the Persuadee experiences a sudden drop in agreement.
We provide two full conversations from PMIYC. Figure 11 illustrates an ideal conversation where both the Persuader and Persuadee adhere to their roles: the Persuader consistently advocates for the claim, while the Persuadee critically evaluates the presented arguments. In contrast, Figure 12 demonstrates the phenomena described in Section 5.3, where the Persuader deviates from their role by advocating for the opposing side, and the Persuadee expresses agreement with the new commonly shared idea rather than the claim itself.
Appendix E PMIYC Prompts
This section presents all the prompts used in PMIYC and their respective use cases. Figure 13 shows the system prompt used for the Persuader in the subjective domain. The Persuader is tasked with persuading the Persuadee to fully agree with the given claim. We employ specific tags to extract agent responses and scores. Similarly, Figure 14 provides the system prompt for the Persuadee in the subjective domain, instructing the model to be reasonable in either accepting or rejecting a claim. The key difference between this prompt and the Persuader’s is that the Persuadee is not inherently biased toward agreement but rather makes a considered decision. To initiate the conversation, the Persuadee receives the initial prompt in Figure 17. We observed that reminding the Persuadee to use tags after each turn improved the success rate of conversation generation. Therefore, we incorporated a reminder prompt, as shown in Figure 16. Finally, at the end of the conversation, the Persuadee receives a final decision prompt (Figure 15). This prompt informs the Persuadee that the conversation has ended, explicitly restates the claim, and requests a final decision. The prompts for the misinformation domain is same with the exception of models being provided a question-answer pair instead of a single claim.
Figure 13: Complete system prompt for the Persuader in the subjective domain.
Figure 14: Complete system prompt for the Persuadee in the subjective domain.
Figure 15: Final decision prompt.
Figure 16: Reminder prompt.
Figure 17: Initial prompt.
Appendix F Experimental Specifications
F.1 Model Specifications
The specific model versions used in our experiments are GPT-4o-mini (2024-07-18), GPT-4o (2024-11-20), Claude 3 Opus (2024-02-29), Claude 2 (Legacy version), and Claude 3 Haiku (2024-03-07). We served the Llama models using LangChain (LangChain, 2023), while GPT-4o and GPT-4o-mini were accessed via Azure OpenAI. Claude models were served using Anthropic’s API.
F.2 Hardware Specifications
The Llama-3.1-8B-Instruct model was hosted on a single NVIDIA A40 GPU, while the Llama-3.3-70B-Instruct model was deployed on two NVIDIA H100 GPUs. For model training in Section 4.1.2 we use four NVIDIA H100 GPUs.
Appendix G License
All code and data used in this project are released under the MIT License. The code is adapted from NegotiationArena, which is also licensed under the MIT License. The Persuasion dataset by Durmus et al. (2024) is licensed under Creative Commons Attribution-NonCommercial-ShareAlike 4.0 (CC BY-NC-SA 4.0), while the Perspectrum dataset by Chen et al. (2019) is licensed under Creative Commons Attribution-ShareAlike (CC BY-SA). The TruthfulQA dataset Lin et al. (2022) is under the Apache License 2.0. The models Llama-3.1-8B-Instruct and Llama-3.3-70B-Instruct are licensed under the Llama 3.1 Community License Agreement and Llama 3.3 Community License Agreement, respectively. Additionally, we utilized GPT-4o and GPT-4o-mini (OpenAI), as well as Claude 3 Opus, Claude 3 Haiku, and Claude 2 (Anthropic), which are proprietary models available under their respective terms of service. Our use of existing artifacts is consistent with their intended use. The artifacts are all in English, and do not contain data with personally identifiable information.